544 mol of sulfuric acid was reacted at constant pressure with 707 mol of sodium hydroxide. This reaction, initially at 25 ºC was found to release 41000 kJ of heat.

Calculate the molar enthalpy change for the following reaction:

H2SO4(aq) + 2NaOH(aq) → 2H2O(aq) + Na2SO4(aq)
Answer #1

From reaction,

1 mol of H2SO4 needs 2 mol of NaOH for complete reaction

So,

544 mol of H2SO4 will need 2*544 = 1088 mol of NaOH

But we have only 707 mol of NaOH.

So, NaOH is limiting reagent.

We will use NaOH is further calculations.

When 707 mol of NaOH reacts, heat released = 41000 KJ

So,

When 2 mol of NaOH reacts, heat released = 41000*2/707 = 116 KJ

In the reaction 2 mol of NaOH are reacting

Since this is heat released, sign would be negative.

Answer: -116 KJ/mol.rxn